OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18

OpenStudy (anonymous):

@johnweldon1993 can you help me

OpenStudy (anonymous):

what are the answer choices

OpenStudy (anonymous):

12x + 4y = -12 15x = -30 8x + 4y = -4 14x = -10 6x - 8y = -36 -6y = -42 6x - y = -15 3y = 9

OpenStudy (anonymous):

@aremmm